OpenStudy (adi3):

suppose you have a = -0.55 b = 12 c = 15 Then which one of the quadratic functions will you use?

OpenStudy (alekos):

need more info to answer that one

OpenStudy (shadowlegendx):

^

OpenStudy (gabebae):

Nah, you don't need more information. You just use the quadratic formula

OpenStudy (alekos):

then go right ahead pal

OpenStudy (adi3):

we can olny use quadratic functions

OpenStudy (adi3):

not quadratic equation

OpenStudy (gabebae):

You need to graph?

OpenStudy (alekos):

are you referring to ax^2+bx+c ?

OpenStudy (adi3):

any function

OpenStudy (alekos):

send me a screenshot of the problem. you're not making any sense

OpenStudy (adi3):

that is the question

OpenStudy (adi3):

we can use any quadratic function to solve the answer

OpenStudy (alekos):

OK, good luck then

OpenStudy (adi3):

lets use ax^2 + bx + c

OpenStudy (adi3):

-0.55x^2 + 12x + 15 then how will you solve it?

OpenStudy (trojanpoem):

Use the quadratic formula -b +- sqrt(b^2 - 4ac)/ 2a

OpenStudy (trojanpoem):

I assume you're trying to find the roots of a function whose coefficients are as mentioned.

OpenStudy (alekos):

you can also try desmos

OpenStudy (adi3):

but it is quadratic functions

Directrix (directrix):

This is the quadratic function: y = -0.55x^2 + 12x + 15. Its graph is a parabola. If you want to find where the graph of the quadratic function intersects the x-axis, you can use the Quadratic Formula for that. The y-value of the points on the x-axis is 0, so set y equal to 0. 0 = -0.55x^2 + 12x + 15 The two x values you get from solving that equation are the x-intercepts of the quadratic function. They are also known as roots or zeroes of the function.

OpenStudy (adi3):

yeah but how to convert from standard form to vertex form

OpenStudy (adi3):

@Directrix

OpenStudy (adi3):

@imqwerty

imqwerty (imqwerty):

what do u knw about the vertex form

OpenStudy (adi3):

a(x-h)^2 + k

OpenStudy (adi3):

sorry

OpenStudy (adi3):

how to covert it to (x, ) (x, )

OpenStudy (adi3):

like when we solve the standard form then we get some thing like this (x, ) (x, ). that what i want

OpenStudy (adi3):

qwerty

imqwerty (imqwerty):

LOL ok XD

imqwerty (imqwerty):

do u want to find the vertex cordinates?

OpenStudy (adi3):

the x intercepts

imqwerty (imqwerty):

the x intercepts are the roots of this equation and u want to find the roots by converting it to vertex form?

OpenStudy (adi3):

morel ike intercept form

imqwerty (imqwerty):

what do u knw about intercept form

OpenStudy (adi3):

a(x-p)(x-q)

imqwerty (imqwerty):

ok so lets convert it into this form we have ax^2 +bx+c=0 we need to convert it into a(x-p)(x-q) ax^2 +x[-p-q]+pq=0 compare it with ax^2 +bx+c=0 we see that b=-p-q c=pq u have b and c u can find p and q then just substitute pa nd q in ur equation :D

OpenStudy (adi3):

i am in 10th man

imqwerty (imqwerty):

so

OpenStudy (adi3):

how about just solve for standard form, we did not learn that yet

OpenStudy (adi3):

yaar

imqwerty (imqwerty):

(:

OpenStudy (adi3):

do you know how to solve the standard form

OpenStudy (adi3):

QWERTY

imqwerty (imqwerty):

well what u call as standard form is basically this- ax^2 +bx+c=0 u can solve it in diff ways the best way is to use the formula like directrix did

OpenStudy (adi3):

remember ac = -0.75, b = 12

OpenStudy (adi3):

QWERTY

OpenStudy (adi3):

dudeeeee

OpenStudy (adi3):

you know what screw this question

imqwerty (imqwerty):

where are u stuck?

OpenStudy (alekos):

Adi3 , If only you told us clearly what you wanted in the first place I could have helped you and you wouldn't have gone through all this grief. Now you're just frustrated and you don't even have an answer

OpenStudy (adi3):

yeah

OpenStudy (alekos):

so, are you after the vertex form of the quadratic equation?

OpenStudy (alekos):

or do you want the x intercepts?

OpenStudy (adi3):

x intercepts